In the context of the Shapiro-Stiglitz model of efficiency wages, what’s the value of effort if
monitoring is costly, the probability that she leaves for another job is b=0.16, the probability that he
gets caught shirking and thus fired is q=0.04 and the efficiency wage in equilibrium is 3,800? (Recall
that the value of working is (W-E)/b and the value of shirking is W/ (b+q))
Answer #1

As long as value of working is higher than value of shirking individual can exert efforts

Therefore we need to find value of E such that

Value of working <=Value of shirking

(W-E)/b<=W/(b+q)

Wb+Wq-E(b+q)<=Wb

Wq-E(b+q)<=0

Wq<=E(b+q)

3800*0.04<=E(0.16+0.04)

152<=E(0.2)

E>=760 therefore if Effort level is above 760 units then individual will shirk

Therefore value of E should be less than 760 units to avoid monitoring.